# NOLO: Navigate Only Look Once

### v1

@@ B.3 Semantic Actions
Appendix C Experiment Details
C.1 Dataset Division
@@
S9hNv5qa7GM XcA2TqTSSAj PX4nDJXEHrG 5LpN3gDmAk7 WYY7iVyf5p8
5q7pvUzZiYa VFuaQ6m2Qom EDJbREhghzL rPc6DW4iMge 2n8kARJN3HM
HxpKQynjfin PuKPg4mmafe 759xd9YjKW5 aayBHfsNo7d jtcxE69GiFV
B6ByNegPMKs p5wJjkQkbXX D7N2EKCX4Sj 1LXtFkjw3qL YVUC4YcDtcY
YmJkqBEsHnH 7y3sRwLe3Va r47D5H71a5s vyrNrziPKCB gxdoqLR6rwA
1pXnuDYAj8r VLzqgDo317F kEZ7cmS4wCh sT4fr6TAbpF ARNzJeq3xxb
Vvot9Ly1tCj cV4RVeZvu5T VVfe2KiqLaN jh4fc5c5qoQ 2t7WUuJeko7
D7G3Y4RVNrH uNb9QFRL6hY V2XKFyX4ASd e9zR4mvMWw7 RPmz2sHmrrY
E9uDoFAP3SH 17DRP5sb8fy JeFG25nYj2p mJXqzFtmKg4 YFuZgdQ5vWj
ac26ZMwG7aT JF19kD82Mey Pm6F8kyY3z2 sKLMLpTHeUy wc2JMjhGNzB
dhjEzFoUFzH 82sE5b5pLXE gZ6f7yhEvPG 8WUmhLawc2A 5ZKStnWn8Zo
pRbA3pwrgk9 29hnd4uzFmX GdvgFV5R1Z5 ULsKaCPVFJR q9vSo1VnCiC
Uxmj2M2itWa qoiz87JEwZ2 s8pcmisQ38h ur6pFq6Qu1A fzynW3qQPVF
ZMojNkEp431 r1Q1Z4BcV1o b8cTxDM8gDG i5noydFURQK JmbYfDe2QKZ
@@
2azQ1b91cZZ x8F5xyUWy9e QUCTc6BB5sX Vt2qJdWjCF2 gTV8FGcVJC9
EU6Fwq7SyZv zsNo4HB9uLZ X7HyMhZNoso yqstnuAEVhm SN83YJsR3w2
pLe4wQe7qrG 8194nk5LbLH Z6MFQCViBuw pa4otMbVnkk VzqfbhrpDEA
TbHJrupSAjP oLBMNvg9in8 rqfALeAoiTq UwV83HsGsw3 gYvKGZ5eRqb
C.2 LMM Baselines
@@
We then utilize regular expressions to extract actions from the received language responses. In case of invalid output, we attempt up to three retries. If all attempts fail, a random action is then selected from the action space as output.
C.3 Hyperparameters
@@
Type of GPUs A100, or Nvidia RTX 4090 Ti
C.4 Computation Resources
@@
Appendix D Evaluation Results
@@
D.1 RoboTHOR Video Navigation Performance
Table 4: NOLO Performance on RoboTHOR Testing Set
@@ D.1 RoboTHOR Video Navigation Performance
FloorPlan_Train10_5 77.45$\pm$ 4.36 28.21$\pm$ 1.58 216.35$\pm$ 14.21 1.68$\pm$ 0.13
FloorPlan_Train11_5 76.44$\pm$ 3.82 36.19$\pm$ 2.32 218.66$\pm$ 13.06 1.62$\pm$ 0.16
FloorPlan_Train12_5 76.27$\pm$ 3.54 31.89$\pm$ 0.94 228.36$\pm$ 4.96 1.50$\pm$ 0.09
FloorPlan_Train1_5 59.72$\pm$ 2.39 33.15$\pm$ 1.72 260.69$\pm$ 9.97 2.06$\pm$ 0.05
FloorPlan_Train2_5 72.54$\pm$ 0.81 28.83$\pm$ 1.45 231.64$\pm$ 4.56 1.76$\pm$ 0.11
FloorPlan_Train3_5 60.20$\pm$ 1.21 23.29$\pm$ 0.21 289.70$\pm$ 6.56 2.34$\pm$ 0.04
FloorPlan_Train4_5 77.42$\pm$ 2.74 27.64$\pm$ 1.66 216.64$\pm$ 6.83 1.71$\pm$ 0.13
FloorPlan_Train5_5 77.58$\pm$ 2.27 32.03$\pm$ 1.45 227.46$\pm$ 4.43 1.54$\pm$ 0.09
FloorPlan_Train6_5 74.44$\pm$ 4.37 30.13$\pm$ 1.04 217.56$\pm$ 6.36 1.90$\pm$ 0.21
FloorPlan_Train7_5 80.00$\pm$ 1.77 22.41$\pm$ 1.20 239.99$\pm$ 18.16 1.39$\pm$ 0.11
FloorPlan_Train8_5 66.25$\pm$ 8.84 27.60$\pm$ 3.87 249.83$\pm$ 31.73 2.07$\pm$ 0.35
FloorPlan_Train9_5 64.67$\pm$ 3.40 29.72$\pm$ 1.78 270.19$\pm$ 12.79 1.95$\pm$ 0.11
@@
FloorPlan_Val1_1 70.83$\pm$ 3.06 30.72$\pm$ 0.78 241.85$\pm$ 11.56 1.74$\pm$ 0.17
FloorPlan_Val1_2 60.45$\pm$ 3.04 23.57$\pm$ 1.53 285.97$\pm$ 5.54 2.31$\pm$ 0.07
FloorPlan_Val1_3 82.98$\pm$ 2.37 36.90$\pm$ 2.38 178.16$\pm$ 11.44 1.41$\pm$ 0.11
FloorPlan_Val1_4 66.85$\pm$ 2.05 26.25$\pm$ 2.42 277.40$\pm$ 15.40 1.85$\pm$ 0.19
FloorPlan_Val1_5 71.81$\pm$ 2.27 34.42$\pm$ 0.91 238.96$\pm$ 5.10 1.87$\pm$ 0.12
FloorPlan_Val2_1 56.97$\pm$ 2.61 17.31$\pm$ 0.98 326.74$\pm$ 9.13 2.30$\pm$ 0.20
FloorPlan_Val2_2 70.56$\pm$ 5.93 31.03$\pm$ 1.34 241.85$\pm$ 24.72 1.92$\pm$ 0.17
FloorPlan_Val2_3 76.41$\pm$ 3.16 33.33$\pm$ 2.68 216.07$\pm$ 8.19 1.49$\pm$ 0.11
FloorPlan_Val2_4 64.36$\pm$ 3.57 25.24$\pm$ 0.84 263.61$\pm$ 15.84 2.04$\pm$ 0.15
FloorPlan_Val2_5 63.33$\pm$ 3.02 24.99$\pm$ 0.36 284.43$\pm$ 5.26 2.44$\pm$ 0.12
FloorPlan_Val3_1 72.86$\pm$ 1.03 26.81$\pm$ 2.75 240.61$\pm$ 8.40 1.90$\pm$ 0.07
FloorPlan_Val3_2 79.31$\pm$ 4.62 27.97$\pm$ 1.41 213.71$\pm$ 14.15 1.60$\pm$ 0.23
FloorPlan_Val3_3 81.48$\pm$ 3.02 30.05$\pm$ 1.48 213.88$\pm$ 10.77 1.46$\pm$ 0.08
FloorPlan_Val3_4 63.81$\pm$ 5.10 18.56$\pm$ 0.80 283.22$\pm$ 10.86 2.11$\pm$ 0.27
FloorPlan_Val3_5 75.14$\pm$ 0.20 28.88$\pm$ 2.25 220.18$\pm$ 9.62 1.65$\pm$ 0.04
@@
Table 5: Random Performance on RoboTHOR Testing Set
@@
FloorPlan_Train10_5 27.15$\pm$ 0.50 16.55$\pm$ 0.85 371.96$\pm$ 3.07 2.84$\pm$ 0.00
FloorPlan_Train11_5 22.83$\pm$ 4.50 11.98$\pm$ 2.67 406.66$\pm$ 15.35 3.24$\pm$ 0.22
FloorPlan_Train12_5 20.68$\pm$ 1.26 11.93$\pm$ 0.51 413.58$\pm$ 4.25 3.02$\pm$ 0.04
FloorPlan_Train1_5 22.00$\pm$ 2.00 13.41$\pm$ 1.47 396.63$\pm$ 0.20 3.08$\pm$ 0.09
FloorPlan_Train2_5 32.83$\pm$ 0.93 14.00$\pm$ 0.57 369.18$\pm$ 2.40 2.54$\pm$ 0.06
FloorPlan_Train3_5 21.26$\pm$ 6.38 10.25$\pm$ 0.81 421.76$\pm$ 1.69 3.36$\pm$ 0.09
FloorPlan_Train4_5 21.32$\pm$ 2.23 9.23$\pm$ 1.62 415.40$\pm$ 9.71 3.08$\pm$ 0.06
FloorPlan_Train5_5 34.95$\pm$ 4.14 16.82$\pm$ 1.57 373.65$\pm$ 12.36 2.55$\pm$ 0.16
FloorPlan_Train6_5 33.67$\pm$ 2.00 17.65$\pm$ 1.43 357.99$\pm$ 0.93 2.76$\pm$ 0.05
FloorPlan_Train7_5 30.75$\pm$ 0.75 12.89$\pm$ 0.42 391.34$\pm$ 6.28 2.54$\pm$ 0.12
FloorPlan_Train8_5 27.00$\pm$ 1.75 11.85$\pm$ 1.68 407.04$\pm$ 8.17 3.16$\pm$ 0.17
FloorPlan_Train9_5 17.50$\pm$ 1.50 12.73$\pm$ 0.25 412.05$\pm$ 6.07 2.96$\pm$ 0.02
@@
FloorPlan_Val1_1 14.50$\pm$ 2.50 9.64$\pm$ 0.75 436.59$\pm$ 3.65 3.13$\pm$ 0.22
FloorPlan_Val1_2 14.50$\pm$ 4.50 8.63$\pm$ 1.24 433.62$\pm$ 8.48 3.25$\pm$ 0.32
FloorPlan_Val1_3 31.61$\pm$ 6.87 16.30$\pm$ 0.77 382.54$\pm$ 14.30 2.77$\pm$ 0.13
FloorPlan_Val1_4 12.28$\pm$ 3.28 8.49$\pm$ 0.73 436.72$\pm$ 9.52 3.18$\pm$ 0.11
FloorPlan_Val1_5 16.17$\pm$ 1.33 10.29$\pm$ 0.50 424.95$\pm$ 9.29 3.23$\pm$ 0.12
FloorPlan_Val2_1 5.41$\pm$ 3.68 4.71$\pm$ 1.47 464.95$\pm$ 11.81 3.17$\pm$ 0.03
FloorPlan_Val2_2 25.61$\pm$ 0.06 13.89$\pm$ 0.33 395.96$\pm$ 1.47 3.17$\pm$ 0.12
FloorPlan_Val2_3 31.04$\pm$ 2.04 17.05$\pm$ 0.02 377.52$\pm$ 5.43 2.17$\pm$ 0.06
FloorPlan_Val2_4 23.73$\pm$ 3.96 14.14$\pm$ 0.76 398.90$\pm$ 3.78 2.50$\pm$ 0.10
FloorPlan_Val2_5 18.20$\pm$ 1.37 11.74$\pm$ 0.15 417.68$\pm$ 0.58 3.25$\pm$ 0.08
FloorPlan_Val3_1 30.45$\pm$ 0.50 13.00$\pm$ 0.84 381.07$\pm$ 4.86 2.55$\pm$ 0.01
FloorPlan_Val3_2 34.08$\pm$ 1.75 12.83$\pm$ 0.82 368.68$\pm$ 1.57 2.79$\pm$ 0.06
FloorPlan_Val3_3 28.39$\pm$ 0.87 11.83$\pm$ 0.51 394.93$\pm$ 4.61 2.84$\pm$ 0.05
FloorPlan_Val3_4 25.21$\pm$ 3.83 10.30$\pm$ 1.26 397.19$\pm$ 9.02 3.01$\pm$ 0.15
FloorPlan_Val3_5 35.33$\pm$ 3.00 17.42$\pm$ 1.22 376.74$\pm$ 3.21 2.34$\pm$ 0.07
@@
Table 6: GPT-4o Performance on RoboTHOR Testing Set
@@
FloorPlan_Train10_5 35.29 19.09 358.81 3.00
FloorPlan_Train11_5 36.00 18.32 363.55 3.15
FloorPlan_Train12_5 35.29 20.42 359.38 3.17
FloorPlan_Train1_5 33.33 21.14 362.02 2.78
FloorPlan_Train2_5 28.57 13.10 367.66 2.76
FloorPlan_Train3_5 20.00 10.86 377.68 3.30
FloorPlan_Train4_5 37.27 18.58 357.89 3.12
FloorPlan_Train5_5 32.73 18.66 363.37 2.95
FloorPlan_Train6_5 41.67 21.92 351.87 3.24
FloorPlan_Train7_5 21.25 8.14 374.45 3.10
FloorPlan_Train8_5 36.25 19.38 356.12 3.07
FloorPlan_Train9_5 26.00 16.57 366.02 2.72
@@
FloorPlan_Val1_1 24.00 13.84 374.19 3.45
FloorPlan_Val1_2 21.82 11.66 373.75 3.22
FloorPlan_Val1_3 46.32 26.44 342.01 2.53
FloorPlan_Val1_4 15.56 8.23 385.98 3.48
FloorPlan_Val1_5 16.59 5.07 390.28 3.10
FloorPlan_Val2_1 28.43 17.75 365.67 3.31
FloorPlan_Val2_2 48.95 23.77 346.67 2.09
FloorPlan_Val2_3 58.52 30.56 329.43 1.82
FloorPlan_Val2_4 35.23 20.31 360.92 2.09
FloorPlan_Val2_5 26.41 14.60 370.46 3.04
FloorPlan_Val3_1 7.48 2.94 391.37 4.18
FloorPlan_Val3_2 41.67 21.22 349.08 2.99
FloorPlan_Val3_3 31.11 13.08 369.33 2.73
FloorPlan_Val3_4 25.71 11.31 373.28 3.22
FloorPlan_Val3_5 47.50 26.64 346.99 2.09
@@
Table 7: Video-LLaVA Performance on RoboTHOR Testing Set
@@
FloorPlan_Train10_5 25.88 13.74 370.67 2.86
FloorPlan_Train11_5 22.67 8.79 380.99 3.58
FloorPlan_Train12_5 17.65 8.67 379.87 3.30
FloorPlan_Train1_5 20.00 7.78 383.98 2.77
FloorPlan_Train2_5 20.95 8.18 382.31 2.67
FloorPlan_Train3_5 14.12 5.89 388.76 3.84
FloorPlan_Train4_5 19.09 6.39 385.62 3.31
FloorPlan_Train5_5 18.18 7.55 385.68 3.04
FloorPlan_Train6_5 23.84 8.69 378.77 2.78
FloorPlan_Train7_5 19.86 7.67 382.04 3.01
FloorPlan_Train8_5 15.79 7.32 384.94 2.84
FloorPlan_Train9_5 22.83 9.11 379.34 2.98
@@
FloorPlan_Val1_1 25.92 10.42 374.65 3.01
FloorPlan_Val1_2 13.64 5.17 389.82 3.44
FloorPlan_Val1_3 10.92 1.66 396.36 3.65
FloorPlan_Val1_4 20.47 8.40 382.13 3.13
FloorPlan_Val1_5 9.17 3.28 395.26 3.63
FloorPlan_Val2_1 14.21 6.74 388.84 3.61
FloorPlan_Val2_2 22.12 8.90 377.86 3.30
FloorPlan_Val2_3 26.38 12.39 372.44 3.13
FloorPlan_Val2_4 13.53 4.34 391.22 4.01
FloorPlan_Val2_5 20.85 6.12 387.35 3.21
FloorPlan_Val3_1 20.00 6.59 381.67 2.93
FloorPlan_Val3_2 17.46 6.13 387.79 2.94
FloorPlan_Val3_3 16.29 5.24 388.70 2.91
FloorPlan_Val3_4 14.36 5.76 388.94 3.55
FloorPlan_Val3_5 17.38 7.51 382.77 3.34
@@
Table 8: NOLO-C Performance on RoboTHOR Testing Set

### v2

@@ 6.3 Semantic Actions
6.4 Model Architecture
[…]
7 Experiment Details
7.1 Dataset Division
@@
S9hNv5qa7GM XcA2TqTSSAj PX4nDJXEHrG
5LpN3gDmAk7 WYY7iVyf5p8 5q7pvUzZiYa
VFuaQ6m2Qom EDJbREhghzL rPc6DW4iMge
2n8kARJN3HM HxpKQynjfin PuKPg4mmafe
759xd9YjKW5 aayBHfsNo7d jtcxE69GiFV
B6ByNegPMKs p5wJjkQkbXX D7N2EKCX4Sj
1LXtFkjw3qL YVUC4YcDtcY YmJkqBEsHnH
7y3sRwLe3Va r47D5H71a5s vyrNrziPKCB
gxdoqLR6rwA 1pXnuDYAj8r VLzqgDo317F
kEZ7cmS4wCh sT4fr6TAbpF ARNzJeq3xxb
Vvot9Ly1tCj cV4RVeZvu5T VVfe2KiqLaN
jh4fc5c5qoQ 2t7WUuJeko7 D7G3Y4RVNrH
uNb9QFRL6hY V2XKFyX4ASd e9zR4mvMWw7
RPmz2sHmrrY E9uDoFAP3SH 17DRP5sb8fy
JeFG25nYj2p mJXqzFtmKg4 YFuZgdQ5vWj
ac26ZMwG7aT JF19kD82Mey Pm6F8kyY3z2
sKLMLpTHeUy wc2JMjhGNzB dhjEzFoUFzH
82sE5b5pLXE gZ6f7yhEvPG 8WUmhLawc2A
5ZKStnWn8Zo pRbA3pwrgk9 29hnd4uzFmX
GdvgFV5R1Z5 ULsKaCPVFJR q9vSo1VnCiC
Uxmj2M2itWa qoiz87JEwZ2 s8pcmisQ38h
ur6pFq6Qu1A fzynW3qQPVF ZMojNkEp431
r1Q1Z4BcV1o b8cTxDM8gDG i5noydFURQK
JmbYfDe2QKZ
@@
2azQ1b91cZZ x8F5xyUWy9e QUCTc6BB5sX
Vt2qJdWjCF2 gTV8FGcVJC9 EU6Fwq7SyZv
zsNo4HB9uLZ X7HyMhZNoso yqstnuAEVhm
SN83YJsR3w2 pLe4wQe7qrG 8194nk5LbLH
Z6MFQCViBuw pa4otMbVnkk VzqfbhrpDEA
TbHJrupSAjP oLBMNvg9in8 rqfALeAoiTq
UwV83HsGsw3 gYvKGZ5eRqb
7.2 LMM Baselines
@@
We then utilize regular expressions to extract actions from the received language responses. In case of invalid output, we attempt up to three retries, and a random action is selected as output if all attempts fail.
7.3 Hyperparameters
@@
Type of GPUs A100 & Nvidia RTX 4090 Ti
7.4 Computation Resources
@@
8 Additional Visualizations
Figure 8: Comparision of context embedding sequences of a sampled video from RoboTHOR between NOLO and NOLO-T.
Sec. 8 compares the t-SNE visual embedding sequence of a context video in RoboTHOR between NOLO and NOLO-T.
9 Evaluation Results
@@
9.1 RoboTHOR Evaluation Performance
We also examined the impact of NOLO compared to LMM baselines and visual navigation baselines on two key metrics for vision-based navigation tasks: Success Rate (SR) and Success weighted by normalized inverse Path Length (SPL), as illustrated in the figures in Fig. 9 and Fig. 10. Figs. 9(a), 9(b), 10(a) and 10(b) show the performance in the unseen layout testing set. Figs. 9(c), 9(d), 10(c) and 10(d) show the evaluation results in the unseen room testing set, and The detailed performances of NOLO, random policy, GPT-4o, Video-LLaVA, AVDC, VGM, ZSON, NOLO-C, NOLO-T, NOLO(M) in RoboTHOR are reported in Tabs. 5, 6, 7, 8, 9, 10, 11, 12, 13 and 14 respectively.
(a) Unseen Layout-SR
(b) Unseen Layout-SPL
(c) Unseen Room-SR
(d) Unseen Room-SPL
Figure 9: Compare NOLO with random policy and LMM baselines in RoboTHOR seen and unseen rooms.
(a) Unseen Layout-SR
(b) Unseen Layout-SPL
(c) Unseen Room-SR
(d) Unseen Room-SPL
Figure 10: Compare NOLO with visual navigation baselines in RoboTHOR seen and unseen rooms.
Table 5: NOLO performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@ 9.1 RoboTHOR Evaluation Performance
T10_5 77.45$\pm$ 4.36 28.21$\pm$ 1.58 216.35$\pm$ 14.21 1.68$\pm$ 0.13
T11_5 76.44$\pm$ 3.82 36.19$\pm$ 2.32 218.66$\pm$ 13.06 1.62$\pm$ 0.16
T12_5 76.27$\pm$ 3.54 31.89$\pm$ 0.94 228.36$\pm$ 4.96 1.50$\pm$ 0.09
T1_5 59.72$\pm$ 2.39 33.15$\pm$ 1.72 260.69$\pm$ 9.97 2.06$\pm$ 0.05
T2_5 72.54$\pm$ 0.81 28.83$\pm$ 1.45 231.64$\pm$ 4.56 1.76$\pm$ 0.11
T3_5 60.20$\pm$ 1.21 23.29$\pm$ 0.21 289.70$\pm$ 6.56 2.34$\pm$ 0.04
T4_5 77.42$\pm$ 2.74 27.64$\pm$ 1.66 216.64$\pm$ 6.83 1.71$\pm$ 0.13
T5_5 77.58$\pm$ 2.27 32.03$\pm$ 1.45 227.46$\pm$ 4.43 1.54$\pm$ 0.09
T6_5 74.44$\pm$ 4.37 30.13$\pm$ 1.04 217.56$\pm$ 6.36 1.90$\pm$ 0.21
T7_5 80.00$\pm$ 1.77 22.41$\pm$ 1.20 239.99$\pm$ 18.16 1.39$\pm$ 0.11
T8_5 66.25$\pm$ 8.84 27.60$\pm$ 3.87 249.83$\pm$ 31.73 2.07$\pm$ 0.35
T9_5 64.67$\pm$ 3.40 29.72$\pm$ 1.78 270.19$\pm$ 12.79 1.95$\pm$ 0.11
@@
V1_1 70.83$\pm$ 3.06 30.72$\pm$ 0.78 241.85$\pm$ 11.56 1.74$\pm$ 0.17
V1_2 60.45$\pm$ 3.04 23.57$\pm$ 1.53 285.97$\pm$ 5.54 2.31$\pm$ 0.07
V1_3 82.98$\pm$ 2.37 36.90$\pm$ 2.38 178.16$\pm$ 11.44 1.41$\pm$ 0.11
V1_4 66.85$\pm$ 2.05 26.25$\pm$ 2.42 277.40$\pm$ 15.40 1.85$\pm$ 0.19
V1_5 71.81$\pm$ 2.27 34.42$\pm$ 0.91 238.96$\pm$ 5.10 1.87$\pm$ 0.12
V2_1 56.97$\pm$ 2.61 17.31$\pm$ 0.98 326.74$\pm$ 9.13 2.30$\pm$ 0.20
V2_2 70.56$\pm$ 5.93 31.03$\pm$ 1.34 241.85$\pm$ 24.72 1.92$\pm$ 0.17
V2_3 76.41$\pm$ 3.16 33.33$\pm$ 2.68 216.07$\pm$ 8.19 1.49$\pm$ 0.11
V2_4 64.36$\pm$ 3.57 25.24$\pm$ 0.84 263.61$\pm$ 15.84 2.04$\pm$ 0.15
V2_5 63.33$\pm$ 3.02 24.99$\pm$ 0.36 284.43$\pm$ 5.26 2.44$\pm$ 0.12
V3_1 72.86$\pm$ 1.03 26.81$\pm$ 2.75 240.61$\pm$ 8.40 1.90$\pm$ 0.07
V3_2 79.31$\pm$ 4.62 27.97$\pm$ 1.41 213.71$\pm$ 14.15 1.60$\pm$ 0.23
V3_3 81.48$\pm$ 3.02 30.05$\pm$ 1.48 213.88$\pm$ 10.77 1.46$\pm$ 0.08
V3_4 63.81$\pm$ 5.10 18.56$\pm$ 0.80 283.22$\pm$ 10.86 2.11$\pm$ 0.27
V3_5 75.14$\pm$ 0.20 28.88$\pm$ 2.25 220.18$\pm$ 9.62 1.65$\pm$ 0.04
@@
Table 6: Random performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@
T10_5 27.15$\pm$ 0.50 16.55$\pm$ 0.85 371.96$\pm$ 3.07 2.84$\pm$ 0.00
T11_5 22.83$\pm$ 4.50 11.98$\pm$ 2.67 406.66$\pm$ 15.35 3.24$\pm$ 0.22
T12_5 20.68$\pm$ 1.26 11.93$\pm$ 0.51 413.58$\pm$ 4.25 3.02$\pm$ 0.04
T1_5 22.00$\pm$ 2.00 13.41$\pm$ 1.47 396.63$\pm$ 0.20 3.08$\pm$ 0.09
T2_5 32.83$\pm$ 0.93 14.00$\pm$ 0.57 369.18$\pm$ 2.40 2.54$\pm$ 0.06
T3_5 21.26$\pm$ 6.38 10.25$\pm$ 0.81 421.76$\pm$ 1.69 3.36$\pm$ 0.09
T4_5 21.32$\pm$ 2.23 9.23$\pm$ 1.62 415.40$\pm$ 9.71 3.08$\pm$ 0.06
T5_5 34.95$\pm$ 4.14 16.82$\pm$ 1.57 373.65$\pm$ 12.36 2.55$\pm$ 0.16
T6_5 33.67$\pm$ 2.00 17.65$\pm$ 1.43 357.99$\pm$ 0.93 2.76$\pm$ 0.05
T7_5 30.75$\pm$ 0.75 12.89$\pm$ 0.42 391.34$\pm$ 6.28 2.54$\pm$ 0.12
T8_5 27.00$\pm$ 1.75 11.85$\pm$ 1.68 407.04$\pm$ 8.17 3.16$\pm$ 0.17
T9_5 17.50$\pm$ 1.50 12.73$\pm$ 0.25 412.05$\pm$ 6.07 2.96$\pm$ 0.02
@@
V1_1 14.50$\pm$ 2.50 9.64$\pm$ 0.75 436.59$\pm$ 3.65 3.13$\pm$ 0.22
V1_2 14.50$\pm$ 4.50 8.63$\pm$ 1.24 433.62$\pm$ 8.48 3.25$\pm$ 0.32
V1_3 31.61$\pm$ 6.87 16.30$\pm$ 0.77 382.54$\pm$ 14.30 2.77$\pm$ 0.13
V1_4 12.28$\pm$ 3.28 8.49$\pm$ 0.73 436.72$\pm$ 9.52 3.18$\pm$ 0.11
V1_5 16.17$\pm$ 1.33 10.29$\pm$ 0.50 424.95$\pm$ 9.29 3.23$\pm$ 0.12
V2_1 5.41$\pm$ 3.68 4.71$\pm$ 1.47 464.95$\pm$ 11.81 3.17$\pm$ 0.03
V2_2 25.61$\pm$ 0.06 13.89$\pm$ 0.33 395.96$\pm$ 1.47 3.17$\pm$ 0.12
V2_3 31.04$\pm$ 2.04 17.05$\pm$ 0.02 377.52$\pm$ 5.43 2.17$\pm$ 0.06
V2_4 23.73$\pm$ 3.96 14.14$\pm$ 0.76 398.90$\pm$ 3.78 2.50$\pm$ 0.10
V2_5 18.20$\pm$ 1.37 11.74$\pm$ 0.15 417.68$\pm$ 0.58 3.25$\pm$ 0.08
V3_1 30.45$\pm$ 0.50 13.00$\pm$ 0.84 381.07$\pm$ 4.86 2.55$\pm$ 0.01
V3_2 34.08$\pm$ 1.75 12.83$\pm$ 0.82 368.68$\pm$ 1.57 2.79$\pm$ 0.06
V3_3 28.39$\pm$ 0.87 11.83$\pm$ 0.51 394.93$\pm$ 4.61 2.84$\pm$ 0.05
V3_4 25.21$\pm$ 3.83 10.30$\pm$ 1.26 397.19$\pm$ 9.02 3.01$\pm$ 0.15
V3_5 35.33$\pm$ 3.00 17.42$\pm$ 1.22 376.74$\pm$ 3.21 2.34$\pm$ 0.07
@@
Table 7: GPT-4o performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@
T10_5 35.29 19.09 358.81 3.00
T11_5 36.00 18.32 363.55 3.15
T12_5 35.29 20.42 359.38 3.17
T1_5 33.33 21.14 362.02 2.78
T2_5 28.57 13.10 367.66 2.76
T3_5 20.00 10.86 377.68 3.30
T4_5 37.27 18.58 357.89 3.12
T5_5 32.73 18.66 363.37 2.95
T6_5 41.67 21.92 351.87 3.24
T7_5 21.25 8.14 374.45 3.10
T8_5 36.25 19.38 356.12 3.07
T9_5 26.00 16.57 366.02 2.72
@@
V1_1 24.00 13.84 374.19 3.45
V1_2 21.82 11.66 373.75 3.22
V1_3 46.32 26.44 342.01 2.53
V1_4 15.56 8.23 385.98 3.48
V1_5 16.59 5.07 390.28 3.10
V2_1 28.43 17.75 365.67 3.31
V2_2 48.95 23.77 346.67 2.09
V2_3 58.52 30.56 329.43 1.82
V2_4 35.23 20.31 360.92 2.09
V2_5 26.41 14.60 370.46 3.04
V3_1 7.48 2.94 391.37 4.18
V3_2 41.67 21.22 349.08 2.99
V3_3 31.11 13.08 369.33 2.73
V3_4 25.71 11.31 373.28 3.22
V3_5 47.50 26.64 346.99 2.09
@@
Table 8: Video-LLaVA performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@
T10_5 25.88 13.74 370.67 2.86
T11_5 22.67 8.79 380.99 3.58
T12_5 17.65 8.67 379.87 3.30
T1_5 20.00 7.78 383.98 2.77
T2_5 20.95 8.18 382.31 2.67
T3_5 14.12 5.89 388.76 3.84
T4_5 19.09 6.39 385.62 3.31
T5_5 18.18 7.55 385.68 3.04
T6_5 23.84 8.69 378.77 2.78
T7_5 19.86 7.67 382.04 3.01
T8_5 15.79 7.32 384.94 2.84
T9_5 22.83 9.11 379.34 2.98
@@
V1_1 25.92 10.42 374.65 3.01
V1_2 13.64 5.17 389.82 3.44
V1_3 10.92 1.66 396.36 3.65
V1_4 20.47 8.40 382.13 3.13
V1_5 9.17 3.28 395.26 3.63
V2_1 14.21 6.74 388.84 3.61
V2_2 22.12 8.90 377.86 3.30
V2_3 26.38 12.39 372.44 3.13
V2_4 13.53 4.34 391.22 4.01
V2_5 20.85 6.12 387.35 3.21
V3_1 20.00 6.59 381.67 2.93
V3_2 17.46 6.13 387.79 2.94
V3_3 16.29 5.24 388.70 2.91
V3_4 14.36 5.76 388.94 3.55
V3_5 17.38 7.51 382.77 3.34
@@
Table 9: AVDC performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.